# Symmetry transformation group arising from the Laplace–Runge–Lenz vector

### v2

@@ 2.1. Infinitesimal symmetry generators from the constants of motion
From Proposition 1, a one-to-one correspondence exists between constants of motion and time-independent infinitesimal dynamical symmetries for the Kepler problem. This correspondence is given explicitly by the relation (2.10). For notational ease, index notation will be adopted hereafter.
@@
$P^{i}[L^{j}]=\frac{\partial{L^{j}}}{\partial{\dot{r}^{i}}}=\epsilon_{ijk}r^{k},\qquad{\bf X}_{L^{j}}=\epsilon_{ijk}r^{k}\partial_{r^{i}}$ (2.27)
(where indices are raised and lowered by the Euclidean metric $\delta_{ij}$). Notice these expressions are strictly linear in $\dot{r}^{i}=v^{i}$, namely $\dfrac{\partial^{2}E}{\partial\dot{r}^{k}\partial\dot{r}^{i}}=\delta_{ik}$ and $\dfrac{\partial^{2}L^{j}}{\partial\dot{r}^{k}\partial\dot{r}^{i}}=0$ are proportional to the identity matrix. Hence, they represent generators of point symmetries.
For energy, the generator ${\bf X}_{E}$ produces the infinitesimal symmetry transformation
$r^{i}\to r^{i}{}^{*}=r^{i}+\varepsilon\dot{r}^{i}+O(\varepsilon^{2})$ (2.28)
with group parameter $\varepsilon$. The resulting group of symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ is given by
$r^{i}(t){}^{*}$ $=\exp(\varepsilon\hat{\bf X}_{E})r^{i}(t)=\exp\Big(\varepsilon\frac{d}{dt}\Big)r^{i}(t)=r^{i}(t+\epsilon),$ (2.29)
$v^{i}(t){}^{*}$ $=\exp(\varepsilon\hat{\bf X}_{E})v^{i}(t)=\exp\Big(\varepsilon\frac{d}{dt}\Big)v^{i}(t)=v^{i}(t+\epsilon)$
where $\hat{\bf X}_{E}=\dot{r}^{i}(t)\partial_{r^{i}}+\dot{v}^{i}(t)\partial_{v^{i}}$. This is clearly a time-translation symmetry group. In particular, it is equivalent to the point transformation
@@
acting in the coordinate space $(t,r^{i},v^{i})$. Note that the minus sign in this transformation comes from the general relation between a point transformation (2.17) and its generator (2.18) on phase space, which shows that $\tau=-1$ and $\xi^{i}=0$.
For angular momentum, the generator ${\bf X}_{L^{j}}$ produces the infinitesimal symmetry transformation
$r^{i}\to r^{i}{}^{*}=r^{i}+\epsilon^{ijk}\varepsilon_{j}r_{k}+O(\varepsilon^{2})$ (2.31)
with group parameter $\varepsilon^{j}$ (which is a vector). This is clearly an infinitesimal rotation acting on $r^{i}$. Note that its generator (2.18) has $\tau=0$ and $\xi^{i}=\epsilon^{ijk}r_{k}$. The corresponding symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ are given by the rotation group
$r^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{L^{j}})r^{i}(t)=R^{i}{}_{j}(\varepsilon)r^{j}(t),\quad v^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{L^{j}})v^{i}(t)=R^{i}{}_{j}(\varepsilon)v^{j}(t)$ (2.32)
with
$R^{i}{}_{j}(\varepsilon)=\cos|\varepsilon|\ \delta^{i}{}_{j}+(1-\cos|\varepsilon|)\hat{\varepsilon}^{i}\hat{\varepsilon}^{j}+\sin|\varepsilon|\ \hat{\varepsilon}_{k}\epsilon^{ki}{}_{j},\quad|\varepsilon|=\varepsilon^{j}\varepsilon_{j},\quad\hat{\varepsilon}^{i}=|\varepsilon|^{-1}\varepsilon^{i}$ (2.33)
being the familiar rotation operator, where $\hat{\bf X}_{L}=\epsilon_{ijk}r^{k}(t)\partial_{r^{i}}+\epsilon_{ijk}v^{k}(t)\partial_{v^{i}}$. In particular, the rotation angle is $|\varepsilon|$ and the rotation axis is $\hat{\varepsilon}^{i}$, determined by the group parameter $\varepsilon^{i}$. These symmetries acts as a point transformation in the coordinate space $(t,r^{i},v^{i})$:
@@
In contrast to the previous generators, this expression is linear in $\dot{r}^{i}=v^{i}$ but not strictly linear, namely $\dfrac{\partial^{2}A^{j}}{\partial\dot{r}^{k}\partial\dot{r}^{i}}=2\delta_{ik}r_{j}-\delta_{jk}r_{i}-\delta_{ij}r_{k}$ is not proportional to $\delta_{ik}$. As a consequence, it represents the generator of a genuine dynamical symmetry. The resulting dynamical symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ involve the prolongation of this generator (2.35) to phase space. Its components consist of $P^{i}[A^{j}]$ and $\mathcal{D}_{t}P^{i}[A^{j}]$, where
$\mathcal{D}_{t}r^{k}=v^{k},\quad\mathcal{D}_{t}v^{k}=-\kappa|r|^{-1}\hat{r}^{k},\quad|r|=\sqrt{r^{k}r_{k}}.$ (2.36)
This yields
$\hat{\bf X}_{A^{j}}=\big(2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k}\big)\partial_{r^{i}}+\big(v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1})\big)\partial_{v^{i}}.$ (2.37)
Thus, on solutions $(r^{i}(t),v^{i}(t))$, the dynamical symmetry transformations arising from the LRL vector are given by
@@
As pointed out in section 1, the underlying constants of motion $E$, $L^{i}$, $A^{i}$ are not all independent, since the magnitude of $A^{i}$ is a function of $E$ and $L^{i}$. The physically relevant content of $A^{i}$ is its direction:
$\Theta^{i}:=|A|^{-1}A^{i},$ (2.39)
where
$|A|=\sqrt{A^{j}A_{j}}=\sqrt{\kappa^{2}+2EL^{j}L_{j}}.$ (2.40)
This direction vector (2.39) is a constant of motion, and hence it gives rise to an infinitesimal symmetry just like $A^{i}$ does. Applying Proposition 1 yields
$P^{i}[\Theta^{j}]=\frac{\partial{\Theta^{j}}}{\partial{\dot{r}^{i}}}=|A|^{-1}(2\dot{r}^{i}r^{j}-r^{i}\dot{r}^{j}-\delta^{ij}\dot{r}^{k}r_{k})+2|A|^{-3}E(r^{i}\dot{r}^{k}-r^{k}\dot{r}^{i})r_{k}A^{j}$ (2.41)
using $|A|^{2}=\kappa^{2}+2EL^{k}L_{k}$ and $L^{k}L_{k}=r_{k}\dot{r}_{l}(r^{k}\dot{r}^{l}-r^{l}\dot{r}^{k})$. Observe that $r^{i}\dot{r}^{k}-r^{k}\dot{r}^{i}=\epsilon^{ikl}L_{l}$, and thus the infinitesimal symmetry generator is given by
${\bf X}_{\Theta^{j}}=\big(|A|^{-1}(2\dot{r}^{i}r^{j}-r^{i}\dot{r}^{j}-\delta^{ij}\dot{r}^{k}r_{k})+2|A|^{-3}E\epsilon^{ikl}L_{l}A^{j}r_{k}\big)\partial_{r^{i}}.$ (2.42)
@@
$\mathcal{D}_{t}P^{i}[\Theta^{j}]=|A|^{-1}(v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1}))+2|A|^{-3}E\epsilon^{ikl}L_{l}A^{j}v_{k}$ (2.43)
by use of the phase-space equations of motion (2.36). Thus,
$\hat{\bf X}_{\Theta^{j}}$ $=\big(|A|^{-1}(2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k})+2|A|^{-3}E\epsilon^{ikl}L_{l}A^{j}r_{k}\big)\partial_{r^{i}}$ (2.44)
$\qquad+\big(|A|^{-1}(v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1}))+2|A|^{-3}E\epsilon^{ikl}L_{l}A^{j}v_{k}\big)\partial_{v^{i}}$
is the prolonged generator in phase space.
The resulting dynamical symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ are then given by
@@
To conclude the present section, the infinitesimal symmetry generators ${\bf X}_{E}$, ${\bf X}_{L^{j}}$, ${\bf X}_{A^{j}}$, and ${\bf X}_{\Theta^{j}}$ will be written in vector notation, as shown in Table 1.
$\begin{gathered}\text{constant}\\ \text{of motion}\end{gathered}$ symmetry generator $\bf X$}
$E$ ${\vec{v}}\cdot\partial_{{\vec{r}}}$
$\vec{L}$ ${\vec{r}}\times\partial_{{\vec{r}}}$
$\vec{A}$ $(2{\vec{r}}\ {\vec{v}}-{\vec{v}}\ {\vec{r}}-{\vec{r}}\cdot{\vec{v}}\ \mathbf{I})\cdot\partial_{{\vec{r}}}$
$\hat{\Theta}$ $|A|^{-1}\big(2{\vec{r}}{\vec{v}}-{\vec{v}}{\vec{r}}+2E|A|^{-2}\vec{A}({\vec{r}}\times\vec{L})-{\vec{r}}\cdot{\vec{v}}\ \mathbf{I}\big)\cdot\partial_{{\vec{r}}}$
Table 1. Generators of infinitesimal symmetries arising from energy, angular momentum, LRL vector, and the LRL direction vector.
@@ 3. Poisson brackets
The Kepler problem (2.1) has the Hamiltonian formulation
$\dot{{\vec{r}}}=\frac{\partial{H({\vec{r}},{\vec{v}})}}{\partial{{\vec{v}}}},\quad\dot{{\vec{v}}}=-\frac{\partial{H({\vec{r}},{\vec{v}})}}{\partial{{\vec{r}}}}$ (3.1)
where
$H({\vec{r}},{\vec{v}})=\tfrac{1}{2}|{\vec{v}}|^{2}-\kappa|{\vec{r}}|^{-1}$ (3.2)
is the energy (2.2). The associated Poisson bracket is given by
@@
for any functions $F({\vec{r}},{\vec{v}})$ and $G({\vec{r}},{\vec{v}})$ on phase space.
As is well known, the constants of motion — energy, angular momentum , LRL vector — form a Lie algebra under the Poisson bracket. To compute the brackets, it is simplest to work in index notation (cf Section 2), with
$E$ $=\tfrac{1}{2}v^{j}v_{j}-\kappa|r|^{-1},$ (3.4)
$L^{i}$ $=\epsilon^{ijk}r_{j}v_{k},$ (3.5)
$A^{i}$ $=\epsilon^{ijk}v_{j}L_{k}-\kappa|r|^{-1}r^{i}=(v^{j}v_{j}-\kappa|r|^{-1})r^{i}-(r^{j}v_{j})v^{i}$ (3.6)
respectively being the energy (Hamiltonian), angular momentum, and LRL vector.
Then, via using expressions (2.26), (2.27), (2.35) obtained from the infinitesimal symmetries, the Poisson brackets are readily computed to be:
$\{E,L^{i}\}$ $=\frac{\partial{E}}{\partial{r^{k}}}P^{k}[L^{i}]-\frac{\partial{L^{i}}}{\partial{r^{k}}}P^{k}[E]=\kappa|r|^{-3}r_{k}\epsilon^{ilk}r_{l}-\epsilon^{ikj}v_{j}v_{k}=0,$ (3.7)
$\{L^{i},L^{j}\}$ $=\frac{\partial{L^{i}}}{\partial{r^{k}}}P^{k}[L^{j}]-\frac{\partial{L^{j}}}{\partial{r^{k}}}P^{k}[L^{i}]=\epsilon^{li}{}_{k}v_{l}\epsilon^{kjm}r_{m}-\epsilon^{lj}{}_{k}v_{l}\epsilon^{kim}r_{m}=\epsilon^{ijk}L_{k},$ (3.8)
and
$\begin{aligned} \{E,A^{i}\}&=\frac{\partial{E}}{\partial{r^{k}}}P^{k}[A^{i}]-\frac{\partial{A^{i}}}{\partial{r^{k}}}P^{k}[E]\\ &=\kappa|r|^{-3}r_{k}(2v^{k}r^{i}-r^{k}v^{i}-\delta^{ki}v^{l}r_{l})-(\delta^{ik}(v^{j}v_{j}-\kappa|r|^{-1})-v^{i}v^{k}+\kappa|r|^{-3}r^{i}r^{k})v_{k}\\ &=0,\end{aligned}$ (3.9)
$\begin{aligned} \{A^{i},L^{j}\}&=\frac{\partial{A^{i}}}{\partial{r^{k}}}P^{k}[L^{j}]-\frac{\partial{L^{j}}}{\partial{r^{k}}}P^{k}[A^{i}]\\ &=(\delta^{ik}(v^{l}v_{l}-\kappa|r|^{-1})-v^{i}v^{k}+\kappa|r|^{-3}r^{i}r^{k})\epsilon^{kjm}r_{m}-\epsilon^{lj}{}_{k}v_{l}(2v^{k}r^{i}-r^{k}v^{i}-\delta^{ki}v^{m}r_{m})\\ &=\epsilon^{ijk}A_{k},\end{aligned}$ (3.10)
$\begin{aligned} \{A^{i},A^{j}\}&=\frac{\partial{A^{i}}}{\partial{r^{k}}}P^{k}[A^{j}]-\frac{\partial{A^{j}}}{\partial{r^{k}}}P^{k}[A^{i}]\\ &=(\delta^{ik}(v^{l}v_{l}-\kappa|r|^{-1})-v^{i}v^{k}+\kappa|r|^{-3}r^{i}r^{k})(2v_{k}r^{j}-r_{k}v^{j}-\delta_{k}{}^{j}v_{m}r^{m})\\ &\qquad-(\delta^{jk}(v^{l}v_{l}-\kappa|r|^{-1})-v^{j}v^{k}+\kappa|r|^{-3}r^{j}r^{k})(2v_{k}r^{i}-r_{k}v^{i}-\delta_{k}{}^{i}v_{m}r^{m})\\ &=-2E\epsilon^{ijk}L_{k}.\end{aligned}$ (3.11)
@@
which has the structure
$\mathbb{R}\oplus\mathfrak{so}(4)\quad\quad\text{ when $E<0$},$ (3.14a)
$\mathbb{R}\oplus\mathfrak{so}(3,1)\quad\text{ when $E>0$}.$ (3.14b)
A key preliminary result now is that this algebra is isomorphic to the Lie algebra formed by the infinitesimal symmetries arising from Noether’s theorem (in reverse) applied to the constants of the motion. This holds in general (see Ref.[4]) for Lagrangian dynamical systems.
Proposition 3.
The infinitesimal symmetries associated to $E$, $L^{i}$, and $M^{i}$ in the Kepler problem have a commutator structure that is isomorphic to the Poisson bracket structure of these constants of motion. Therefore, these infinitesimal symmetries form a Lie algebra (3.14).
The structure of the Lie algebra (3.14) turns out to be closely tied to the use of $M^{i}$ in place of $A^{i}$. Perhaps unsurprisingly, since $M^{i}$ and $L^{i}$ have the same physical dimensions, the Lie subalgebra that they generate cannot be factored into simpler subalgebras. So it is natural to expect that if $M^{i}$ is replaced by the LRL direction vector (2.39), which is dimensionless, then the Lie subalgebra generated by this vector $\Theta^{i}$ and $L^{i}$ should have a simpler structure.
The brackets involving $\Theta^{i}$ can be computed directly from the previous brackets (3.7)–(3.11) by using the relation $A^{i}=|A|\,\Theta^{i}$ together with the property that the Poisson bracket obeys the Leibniz rule. In particular, for any function $F(r^{k},v^{k})$ on phase space, note that its bracket with $\Theta^{i}$ can be obtained from $\{F,A^{i}\}=|A|\ \{F,\Theta^{i}\}+\Theta^{i}\ \{F,|A|\}$, where $\{F,|A|\}=|A|^{-1}A_{j}\ \{F,A^{j}\}$. Hence, firstly,
$\{E,|A|\}$ $=|A|^{-1}A_{j}\ \{E,A^{j}\}=0,$ (3.15)
$\{L^{i},|A|\}$ $=|A|^{-1}A_{j}\ \{L^{i},A^{j}\}=|A|^{-1}\epsilon^{ijk}A_{j}A_{k}=0,$ (3.16)
$\{A^{i},|A|\}$ $=|A|^{-1}A_{j}\ \{A^{i},A^{j}\}=-2|A|^{-1}E\epsilon^{ijk}A_{j}L_{k};$ (3.17)
@@
$\{E,\Theta^{i}\}=0,$ (3.18)
$\{L^{j},\Theta^{i}\}=|A|^{-1}\epsilon^{jik}A_{k}=\epsilon^{jik}\Theta_{k},$ (3.19)
$\{|A|,\Theta^{i}\}=2|A|^{-1}E\epsilon^{ijk}A_{j}L_{k},$ (3.20)
$\{A^{j},\Theta^{i}\}=2|A|^{-1}E(\epsilon^{jlk}\Theta_{l}\Theta^{i}-\epsilon^{jik})L_{k},$ (3.21)
$\begin{aligned} \{\Theta^{j},\Theta^{i}\}&=-2|A|^{-2}E(\epsilon^{jik}-(\epsilon^{jlk}\Theta^{i}-\epsilon^{ilk}\Theta^{j})\Theta_{l}))L_{k}\\ &=-2|A|^{-2}E\epsilon^{jik}((1-\Theta_{l}\Theta^{l})L_{k}+\Theta^{l}L_{l}\Theta_{k})=0,\end{aligned}$ (3.22)
where the last line has used the vector identity $\epsilon^{jlk}F^{i}-\epsilon^{ilk}F^{j}=\epsilon^{jik}F^{l}-\epsilon^{jil}F^{k}$, as well as the relations $\Theta^{j}L_{j}=0$ and $\Theta^{j}\Theta_{j}=1$.
Therefore, only the bracket (3.19) among $E$, $L^{i}$, $\Theta^{i}$ is non-vanishing. This bracket shows that $L^{j}$ acts as an infinitesimal rotation on the direction vector $\Theta^{i}$:
$\{L^{i},\Theta^{j}\}=\epsilon^{ij}{}_{k}\Theta^{k}.$ (3.23)
@@
The constants of motion $E$, $L^{i}$, and $\Theta^{i}$ generate a 7-dimensional Lie algebra under the Poisson bracket, which has the structure
$\mathbb{R}\oplus\mathfrak{so}(3)\rtimes\mathbb{R}^{3}.$ (3.24)
This algebra is isomorphic to the Lie algebra of the infinitesimal symmetries associated to these constants of motion.

### v3

@@ 2.1. Infinitesimal symmetry generators from the constants of motion
Proposition 1 gives a one-to-one Noether correspondence between constants of motion and time-independent infinitesimal dynamical symmetries for the Kepler problem.
@@
$IP^{i}[L^{j}]=\frac{\partial{L^{j}}}{\partial{\dot{r}^{i}}}=\epsilon_{ijk}r^{k},\qquad{\bf X}_{L^{j}}=\epsilon_{ijk}r^{k}\partial_{r^{i}}$ (2.17)
which have the form of point symmetries (2.12), namely $P^{i}$ is strictly linear in $\dot{r}^{i}=v^{i}$.
The generator ${\bf X}_{E}$ produces the infinitesimal symmetry $r^{i}\to r^{i}{}^{*}=r^{i}+\varepsilon\dot{r}^{i}+O(\varepsilon^{2})$ with group parameter $\varepsilon$, yielding the symmetry transformation group on solutions
$r^{i}(t){}^{*}=\exp(\varepsilon\hat{\bf X}_{E})r^{i}(t)=\exp\Big(\varepsilon\frac{d}{dt}\Big)r^{i}(t)=r^{i}(t+\epsilon)$ (2.18)
and $v^{i}(t){}^{*}=\frac{d}{dt}r^{i}(t){}^{*}=v^{i}(t+\epsilon)$, where $\hat{\bf X}_{E}=\dot{r}^{i}(t)\partial_{r^{i}}+\dot{v}^{i}(t)\partial_{v^{i}}$. This is a time-translation symmetry group, which is equivalent to the point transformation
@@
acting in the coordinate space $(t,r^{i},v^{i})$. Note that the minus sign in the transformation of $t$ comes from the general relation between a point transformation (2.11) and its generator (2.12) on phase space, which shows that $\tau=-1$ and $\xi^{i}=0$.
The angular momentum generator ${\bf X}_{L^{j}}$ produces $r^{i}\to r^{i}{}^{*}=r^{i}+\epsilon^{ijk}\varepsilon_{j}r_{k}+O(\varepsilon^{2})$ with group parameter $\varepsilon^{j}$ (which is a vector). This is manifestly an infinitesimal rotation acting on $r^{i}$. Note that its generator (2.12) has $\tau=0$ and $\xi^{i}=\epsilon^{ijk}r_{k}$. The corresponding symmetry transformations on solutions are given by the rotation group
$r^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{L^{j}})r^{i}(t)=R^{i}{}_{j}(\varepsilon)r^{j}(t)$ (2.20)
and $v^{i}(t){}^{*}=\frac{d}{dt}r^{i}(t){}^{*}=R^{i}{}_{j}(\varepsilon)v^{j}(t)$, where $\hat{\bf X}_{L}=\epsilon_{ijk}r^{k}(t)\partial_{r^{i}}+\epsilon_{ijk}v^{k}(t)\partial_{v^{i}}$, with
$R^{i}{}_{j}(\varepsilon)=\cos(|\varepsilon|)\delta^{i}{}_{j}+(1-\cos(|\varepsilon|))\hat{\varepsilon}^{i}\hat{\varepsilon}^{j}+\sin(|\varepsilon|)\hat{\varepsilon}_{k}\epsilon^{ki}{}_{j},\quad|\varepsilon|=\varepsilon^{j}\varepsilon_{j},\quad\hat{\varepsilon}^{i}=|\varepsilon|^{-1}\varepsilon^{i}$ (2.21)
being the familiar rotation operator. In particular, the rotation angle is $|\varepsilon|$ and the rotation axis is $\hat{\varepsilon}^{i}$, determined by the group parameter $\varepsilon^{i}$. These symmetries acts as a point transformation in the coordinate space $(t,r^{i},v^{i})$:
@@
In contrast to the previous generators (2.16) and (2.17), this expression is not strictly linear in $\dot{r}^{i}=v^{i}$, namely $\frac{\partial{P^{i}[A^{j}]}}{\partial{\dot{r}^{k}}}=2\delta_{ik}r_{j}-\delta_{jk}r_{i}-\delta_{ij}r_{k}$ is not proportional to $\delta_{ik}$. As a consequence, it represents the generator of a genuine dynamical symmetry rather than a point symmetry. The resulting dynamical symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ involve the prolongation of this generator (2.23) to phase space. Its components consist of $P^{i}[A^{j}]$ and $\mathcal{D}_{t}P^{i}[A^{j}]$, where $\mathcal{D}_{t}r^{k}=v^{k}$ and $\mathcal{D}_{t}v^{k}=-\kappa r^{-1}\hat{r}^{k}$, with $r=\sqrt{r^{k}r_{k}}=|{\vec{r}}|$. This yields
$\hat{\bf X}_{A^{j}}=\big(2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k}\big)\partial_{r^{i}}+\big(v^{i}v^{j}-\kappa r^{-3}r^{i}r^{j}-\delta^{ij}(2E+\kappa r^{-1})\big)\partial_{v^{i}}.$ (2.24)
Thus, the dynamical symmetry transformations arising from the LRL vector are given by
@@
As pointed out in section 1, the underlying constants of motion $E$, $L^{i}$, $A^{i}$ are not all independent, since the magnitude of $A^{i}$ is a function of $E$ and $|L|=\sqrt{L^{i}L_{i}}$. The physically relevant content of $A^{i}$ is its direction, which is given by the unit vector
$\Theta^{i}:=|A|^{-1}A^{i}$ (2.26)
where $|A|=\sqrt{A^{j}A_{j}}=\sqrt{\kappa^{2}+2E|L|^{2}}$. This direction vector (2.26) is a constant of motion, and hence it gives rise to an infinitesimal symmetry just like $A^{i}$ does. Proposition 1 yields
$P^{i}[\Theta^{j}]=\frac{\partial{\Theta^{j}}}{\partial{\dot{r}^{i}}}=|A|^{-1}(2\dot{r}^{i}r^{j}-r^{i}\dot{r}^{j}-\delta^{ij}\dot{r}^{k}r_{k})-|A|^{-3}(2E\epsilon^{ikl}L_{k}r_{l}+\tfrac{1}{2}|L|^{2}\dot{r}^{i})A^{j}$ (2.27)
Hence, the infinitesimal symmetry generator is given by
${\bf X}_{\Theta^{j}}=\big(|A|^{-1}(2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k})-|A|^{-3}(2E\epsilon^{ikl}L_{k}r_{l}+|L|^{2}v^{i})A^{j}\big)\partial_{r^{i}}.$ (2.28)
@@
$\mathcal{D}_{t}P^{i}[\Theta^{j}]=|A|^{-1}(v^{i}v^{j}-\kappa r^{-3}r^{i}r^{j}-\delta^{ij}(2E+\kappa r^{-1}))-|A|^{-3}(2E\epsilon^{ikl}L_{k}v_{l}-\kappa|L|^{2}r^{-3}r^{i})A^{j}.$ (2.29)
Thus, the prolonged generator in phase space is
$\hat{\bf X}_{\Theta^{j}}$ $=\big(|A|^{-1}(2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k})-|A|^{-3}(2E\epsilon^{ikl}L_{k}r_{l}+|L|^{2}v^{i})A^{j}\big)\partial_{r^{i}}$ (2.30)
$\qquad+\big(|A|^{-1}(v^{i}v^{j}-\kappa r^{-3}r^{i}r^{j}-\delta^{ij}(2E+\kappa r^{-1}))-|A|^{-3}(2E\epsilon^{ikl}L_{k}v_{l}-\kappa|L|^{2}r^{-3}r^{i})A^{j}\big)\partial_{v^{i}},$
and the resulting dynamical symmetry transformations are given by
@@ 3. Poisson brackets
The Kepler problem (2.1) has the Hamiltonian formulation $\dot{{\vec{r}}}=\frac{\partial{H({\vec{r}},{\vec{v}})}}{\partial{{\vec{v}}}}$, $\dot{{\vec{v}}}=-\frac{\partial{H({\vec{r}},{\vec{v}})}}{\partial{{\vec{r}}}}$ where $H({\vec{r}},{\vec{v}})=\tfrac{1}{2}|{\vec{v}}|^{2}-\kappa|{\vec{r}}|^{-1}$ is the energy (2.2). The Poisson bracket is given by
@@
for any functions $F({\vec{r}},{\vec{v}})$ and $G({\vec{r}},{\vec{v}})$ on phase space. When these functions are constants of motion, note that $\frac{\partial{F}}{\partial{{\vec{v}}}}=\vec{P}[F]$ and $\frac{\partial{G}}{\partial{{\vec{v}}}}=\vec{P}[G]$ are given by the corresponding infinitesimal symmetry generators through Noether’s theorem as stated in Proposition 1.
As is well known, the constants of motion — energy, angular momentum, LRL vector — form a Lie algebra under the Poisson bracket. To compute the brackets, it is simplest to work in index notation (cf Section 2) and use expressions (2.16), (2.17), (2.23). This yields
$\{E,L^{i}\}=0,\quad\{L^{i},L^{j}\}=\epsilon^{ijk}L_{k},$ (3.2)
$\{E,A^{i}\}=0,\quad\{A^{i},L^{j}\}=\epsilon^{ijk}A_{k},\quad\{A^{i},A^{j}\}=-2E\epsilon^{ijk}L_{k}.$ (3.3)
@@
which has the structure $\mathbb{R}\oplus\mathfrak{so}(4)$ when $E<0$, and $\mathbb{R}\oplus\mathfrak{so}(3,1)$ when $E>0$.
Perhaps unsurprisingly, since $M^{i}$ and $L^{i}$ have the same physical dimensions, the rotation algebras $\mathfrak{so}(4)$ and $\mathfrak{so}(3,1)$ that they generate cannot be factored into simpler subalgebras. So it is natural to expect that if $M^{i}$ is replaced by the dimensionless LRL direction vector (2.26), then the Lie algebra generated by $\Theta^{i}$ and $L^{i}$ should have a simpler structure.
The brackets involving $\Theta^{i}$ can be computed directly from the previous brackets (3.2)–(3.3) by using the relation $A^{i}=|A|\,\Theta^{i}$ together with the Leibniz rule. In particular, for any function $F(r^{k},v^{k})$ on phase space, note that its bracket with $\Theta^{i}$ can be obtained from $\{F,A^{i}\}=|A|\ \{F,\Theta^{i}\}+\Theta^{i}\ \{F,|A|\}$, where $\{F,|A|\}=|A|^{-1}A_{j}\ \{F,A^{j}\}$. This gives
$\{E,|A|\}=0,\quad\{L^{i},|A|\}=0,\quad\{A^{i},|A|\}=-2|A|^{-1}E\epsilon^{ijk}A_{j}L_{k};$ (3.6)
@@
$\{E,\Theta^{i}\}=0,\quad\{L^{j},\Theta^{i}\}=\epsilon^{jik}\Theta_{k},$ (3.7)
$\{|A|,\Theta^{i}\}=2E\epsilon^{ijk}\Theta_{j}L_{k},\quad\{A^{j},\Theta^{i}\}=2|A|^{-1}E(\epsilon^{jlk}\Theta_{l}\Theta^{i}-\epsilon^{jik})L_{k},\quad\{\Theta^{j},\Theta^{i}\}=0,$ (3.8)
where the last line has used the vector identity $\epsilon^{jlk}F^{i}-\epsilon^{ilk}F^{j}=\epsilon^{jik}F^{l}-\epsilon^{jil}F^{k}$, as well as the relations $\Theta^{j}L_{j}=0$ and $\Theta^{j}\Theta_{j}=1$. Therefore, only the bracket $\{L^{i},\Theta^{j}\}$ among $E$, $L^{i}$, $\Theta^{i}$ is non-zero. It shows that $L^{j}$ acts as an infinitesimal rotation on the direction vector $\Theta^{i}$.
@@
The constants of motion $E$, $L^{i}$, and $\Theta^{i}$ generate a 7-dimensional Lie algebra under the Poisson bracket, which has the structure $\mathbb{R}\oplus\mathfrak{so}(3)\rtimes\mathbb{R}^{3}$. This algebra is isomorphic to the Lie algebra of the infinitesimal symmetries associated to these constants of motion.